OpenStudy (anonymous):

use the foil method (3a+4b)(5a+b)

OpenStudy (angela210793):

3a*5a+3a*b+4b*5a+4b*b=15a^2+3ab+20ab+4b^2=15a^2+23ab+4b^2

OpenStudy (radar):

First, Outer, Inner, Last = FOIL First: 3a 5a=15a^2 Outer: 3a b = 3ab Inner: 4b 5= 20b last: 4b b= 4b^2 Now combine: \[15a ^{2}+3ab+20b+4b ^{2}\]